# Considerations for the planning, conduct and reporting of clinical trials with interim analyses

### v2

@@ 4.1 Multiple endpoints, arms, or subgroups
Most confirmatory clinical trials aim for rejection of more than one null hypothesis and there may be insufficient information available to do this for all endpoints at the same time. In such cases, it will be beneficial to define endpoint-specific sets of analysis timepoints, e.g., interim and primary analyses for PFS and interim and primary analyses for OS in oncology trials. These definitions at the planning stage subsequently may lead to unique analysis time points for the respective confirmatory analyses at the reporting stage, while the decisive analysis is usually based on a single CCOD.
In a group-sequential design, a type 1 error control strategy may need to be considered for all endpoints with hypothesis testing individually [17, 13]—otherwise an inflation of the family-wise type 1 error may occur. Endpoint specific terminology will also facilitate this aspect.
Multiple hypotheses are also tested in designs with more than two arms with either confirmatory adaptive arm selection at an interim analysis (an example is the GATSBY trial, see Thuss et al. [26]), formal pairwise comparisons between all arms as in the CLL11 trial (see Asikanius et al.[1]), or potentially within a platform trial (see e.g. Howard et al.[16]). In such trials, comparisons may occur at different timepoints, thus requiring a different timing of analyses and hence also different CCODs / SSDs per comparison. In particular, an interim analysis for intervention A vs. C may become a confirmatory analysis, while evidence to claim superiority of B vs. C may only be available at a later interim (and then hence confirmatory) analysis. Hence, at reporting stage terminology such as confirmatory analysis for A vs. B and confirmatory analysis for A vs. C may be used.
@@ 4.2.1 Pipeline information
Pipeline information is a term which is of particular relevance for trials with endpoints that are captured at a specific follow-up time. Examples of the former include the 6 minute walk distance at 26 weeks in Pulmonary Hypertension or the clinical remission rate at 52 weeks in Crohn’s Disease.
Often, and especially at interim analyses, at the time of CCOD not all patients will have reached their targeted follow-up. For example, if an interim analysis is conducted when 60 of originally planned 100 patients have reached Week 26 (primary endpoint timing), while 80 have been enrolled, the data on the 20 patients who have been enrolled, but did not reach their primary endpoint yet, is referred to as pipeline information[14]. This includes data between the CCOD and SSDbut also information beyond the SSD. This data is usually not part of the analysis that is reported for this CCOD. The most common exception are pipeline deaths which are reported as part of the complete assessment of safety. However, it may be that pipeline data are deemed important later and used in the decisive analysis to update estimated effects received at the confirmatory analysis (be it an interim or the primary).
@@ 4.2.3 Over- and underrunning for a time-to-event endpoint at the primary analysis
As one can easily deduct based on the reasons for over- or underrunning at an interim analysis, this can also happen at the primary analysis. In that case, the information fractions are updated according to the actually observed maximum information. Updating the significance levels of all analyses based on these new information fractions using the pre-specified $\alpha$-spending function is not permissible, as it would require the modification to the critical boundaries for analyses which were already conduced at earlier interim analyses. The modification of these critical values and hence of those decisions would harm the validity of the trial. Instead, one uses the $\alpha$ that has actually been spent at earlier interim analyses and recomputes the remaining significance level for the primary analysis based on these and the updated information fractions. Wassmer and Brannath [30] (Chapter 3.3) and Wolbers et al. [31] offer a comprehensive summary of this approach including further references. Obviously, this approach to under- and overruning is only valid as long as the under- or overrunning has occurred at random. Scenarios in which the team actively decides to delay the primary analysis due to missing success at an interim analysis would effectively result in a sample-size re-estimation. This approach would result in type I error inflation, unless proper adjustments are being implemented for correction of critical values or test statistics.
@@
5 Description of a group-sequential design in a trial protocol or statistical analysis plan
To conclude this paper we provide template text that could be included in a trial protocol or statistical analysis plan. Note that this is again a hypothetical example only and choices might differ in practice. The reporting of that trial will be discussed in Section 6 below. Note that, other than in this example, in a real trial it would be expected to justify the assumptions and design choices. As this section, as well as Section 6, may serve as templates we do not use any abbreviations on purpose.
@@ 5 Description of a group-sequential design in a trial protocol or statistical analysis plan
At the design stage, names of analysis are given as in the first column of Table 5. The label of the analysis at the analysis stage depends on the outcome of the trial (see Table 5). If the trial stops at interim analysis 1 or interim analysis 2 for futility, the corresponding analysis is also referred to as futility analysis at the reporting stage. If the trial
@@
6 Description of the results of a group-sequential design
In this section we propose how to report the outcome of a trial that was stopped for futility or efficacy, respectively. To this end, we have simulated trials based on the assumptions of Section 5 and under two scenarios for the treatment effect: Trial 1 with an assumed hazard ratio of 1 and Trial 2 with an assumed hazard ratio of 0.75. We repeatedly simulated new trials until we had a trial that was stopped for fultility at interim analysis 1 (Trial 1) and one which was stopped for efficacy at interim analysis 2 (Trial 2). These simulated trials are used as the basis for suggested reporting in Sections 6.1 and 6.2.
When reporting results of a group-sequential design, or more general adaptive design, we strongly recommend to take inspiration from the CONSORT extension for adaptive designs[6], as applicable.
6.1 Reporting of a trial that was stopped for futility
@@
6.2 Reporting of a trial that was stopped for efficacy
@@
7 Discussion
During the past decades, interim analyses implemented through a group-sequential design have become common, especially for trials with time-to-event endpoints. The statistical theory is well established and straightforward to implement at the design stage. Terminology for such designs is, however, not standard, and to some extent, also misleading. Standardized terminology is important to ensure a common understanding of the trial objectives and scope of different analyses among all stakeholders. Here, as a team of statisticians from pharmaceutical industry, regulatory agencies, and academia, we propose terminology for clinical trials with interim analyses and confirmatory intent, i.e. hypothesis testing. The nomenclature can, however, be used flexibly in different types of clinical trial designs, including group sequential trials with multiple endpoints.
8 Declarations
Availability of data and material
Authors’ contributions

### v3

@@ 4.1 Multiple endpoints, arms, or subgroups
Most confirmatory clinical trials aim for rejection of more than one null hypothesis. Frequently, there may be insufficient information available to do this for all endpoints at the same time. In such cases, it will be beneficial to define endpoint-specific sets of analysis time points, e.g., interim and primary analyses for PFS and interim and primary analyses for OS in oncology trials. These definitions at the planning stage subsequently may lead to unique analysis time points for the respective confirmatory analyses at the reporting stage, while the decisive analysis is usually based on a single CCOD.
In a group-sequential design, a type 1 error control strategy will need to be considered for all endpoints with hypothesis testing individually [20, 16]—otherwise an inflation of the family-wise type 1 error may occur. We recommend to use endpoint specific terminology to facilitate this aspect.
Multiple hypotheses are also tested in designs with more than two arms with either confirmatory adaptive arm selection at an interim analysis (an example is the GATSBY trial, see Thuss et al. [28]), formal pairwise comparisons between all arms as in the CLL11 trial (see Asikanius et al.[1]), or potentially within a platform trial (see e.g. Howard et al.[19]). In such trials, comparisons may occur at different time points, thus requiring a different timing of analyses and hence also different CCODs / SSDs per comparison. In particular, an interim analysis for intervention A vs. C may become a confirmatory analysis, while evidence to claim superiority of B vs. C may only be available at a later interim (and then hence confirmatory) analysis. Hence, at reporting stage terminology such as confirmatory analysis for A vs. B and confirmatory analysis for A vs. C should be used.
@@ 4.2 Pipeline information, overrunning and underrunning
A key motivation for the development of this paper was that there is typically a gap between recruitment of study participants and observing their outcome. If all outcomes were available at the same time for all participants, there might be just one single analysis time point. Different terminology has been used in the past interchangeably to describe data which is accruing but has not been reported yet. This subsection aims to clarify the difference between pipeline information and (random) overrunning and how this possibly impacts trial design.
@@ 4.2.1 Pipeline information
Pipeline information is a term which is of particular relevance for trials with endpoints that are captured at a specific follow-up time after randomization. Examples of the former include the 6 minute walk distance at 26 weeks in Pulmonary Hypertension or the clinical remission rate at 52 weeks in Crohn’s Disease.
Often, and especially at interim analyses, at the time of CCOD not all patients will have reached their targeted follow-up. For example, if an interim analysis is conducted when 60 of originally 100 planned patients have reached Week 26 (primary endpoint timing), while 80 have been enrolled, the data on the 20 patients who have been enrolled, but did not reach their primary endpoint yet, is referred to as pipeline information[17]. This includes data between the CCOD and SSDbut also information beyond the SSD. This data is usually not part of the analysis that is reported for this CCOD. The most common exception are pipeline deaths which are reported as part of the complete assessment of safety. However, it may be that pipeline data are deemed important later and used in the decisive analysis to update estimated effects received at the confirmatory analysis (be it an interim or the primary).
@@ 4.2.3 Over- and underrunning for a time-to-event endpoint at the primary analysis
As one can easily deduce based on the reasons for over- or underrunning at an interim analysis, this can also happen at the primary analysis. In that case, the information fractions are updated according to the actually observed maximum information. Updating the significance levels of all analyses based on these new information fractions using the pre-specified $\alpha$-spending function is not permissible, as it would require the modification to the critical boundaries for analyses which were already conduced at earlier interim analyses. The modification of these critical values and hence of those decisions would harm the validity of the trial. Instead, one uses the $\alpha$ that has actually been spent at earlier interim analyses and recomputes the remaining significance level for the primary analysis based on these and the updated information fractions. Wassmer and Brannath [32] (Chapter 3.3) and Wolbers et al. [33] offer a comprehensive summary of this approach including further references. Obviously, this approach to under- and overruning is only valid as long as the under- or overrunning has occurred at random. Scenarios in which the team actively decides to delay the primary analysis due to missing success at an interim analysis would effectively result in a sample-size re-estimation. This approach would result in type I error inflation, unless proper adjustments are being implemented for correction of critical values or test statistics.
@@
4.3 Stopping a trial
In discussing trial designs across stakeholders, the formulation stopping for efficacy and stopping for futility is often used. Indeed, if the null hypothesis corresponding to the primary endpoint can be rejected at an interim analysis for efficacy within a group-sequential design, then the testing for this hypothesis is completed. In statistical terms, the trial is “stopped”. The result is statistically significant and the trial is formally a “success”.
Similarly, if the trial is declared as futile, no further formal hypothesis testing will be conducted and the trial is considered “negative”, even though it may have generated important insights. Notably, a futility declaration may not necessarily mean that the tested intervention is not efficacious, but could also point towards an inadequate trial design.
From an operational perspective, the hypothesis test of the primary endpoint is not directly linked with how patients are treated and followed in a clinical trial. Indeed, data collection in the trial will typically continue after stopping for efficacy, i.e., rejection of the null hypothesis. In some cases, a trial simply continues to allow patients to still receive treatment and/or to collect more data to comprehensively describe and quantify the treatment effect, e.g., on secondary or safety endpoints[6]. An example is in oncology, where PFS may be the primary and OS the secondary endpoint.
Stopping a trial for efficacy usually impacts the conduct of a trial even if the trial is not stopped from an operational perspective and further data is collected: unblinded summary measures and results might be released; the trial sponsor might be unblinded to individual patient data; investigators might be unblinded; patients might be unblinded and/or be allowed to cross-over to the new therapy. All this has the potential to introduce bias and impact the continued collection of reliable and relevant data. Hence, care should be taken at the planning stage to ensure that results at this stage a sufficiently mature for all important objectives. If collection of further information is essential, measures should be considered to mitigate bias if possible. These measures could include clear communication strategies, firewall procedures to protect the core trial team, and strategies how to continue the trial.
A futility analysis reduces the power of a hypothesis test, and therefore also the probability of a type 1 error. Hence, one might be tempted to adjust the critical value of the hypothesis test such that the significance level is again fully exploited, implying a slight power increase. In that case it is imperative to stick to the futility decision as otherwise the type 1 error would no longer be controlled. As the decision to stop a trial for futility is usually multi-faceted and includes amongst others also a review of the safety data, non-binding futility analyses are generally preferable[14]. As these authors explain, the term “non-binding” applies to both its semantic interpretation and its statistical definition: reaching such a threshold does not rigidly mandate stopping, and final efficacy thresholds are not loosened so that a decision to continue does not inflate the type I error. It is important to clearly keep the two meanings of non-binding apart.
5 Discussion
Standardized terminology is important to ensure a common understanding of the objectives and scope of different analyses in a clinical trial among all stakeholders. In this paper, as a team of statisticians from pharmaceutical industry, regulatory agencies, and academia, we propose terminology for clinical trials with interim analyses and confirmatory intent, i.e. hypothesis testing. The nomenclature can, however, be used flexibly in different types of clinical trial designs, including group sequential trials with multiple endpoints.
6 Declarations
Ethics approval and consent to participate
Not applicable.
Consent for publication
Not applicable.
Availability of data and material
Competing interests
Not applicable.
Funding
Not applicable.
Authors’ contributions
Acknowledgements
Not applicable.
Authors’ information (optional)
Not applicable.
Appendix A Description of a group-sequential design in a trial protocol or statistical analysis plan
To conclude this paper we provide illustrative text that could be included in a trial protocol or statistical analysis plan. Note that this is again a hypothetical example only and choices might differ in practice. The reporting of that trial will be discussed in Section B below. Note that, other than in this example, in a real trial it would be expected to justify the assumptions and design choices. As this section, as well as Section B, may serve as templates we do not use any abbreviations on purpose.
@@ Appendix A Description of a group-sequential design in a trial protocol or statistical analysis plan
At the design stage, names of analysis are given as in the first column of Table A. The label of the analysis at the analysis stage depends on the outcome of the trial (see Table A). If the trial stops at interim analysis 1 or interim analysis 2 for futility, the corresponding analysis is also referred to as futility analysis at the reporting stage. If the trial
@@
Appendix B Description of the results of a group-sequential design
In this section we propose how to report the outcome of a trial that was stopped for futility or efficacy, respectively. To this end, we have simulated trials based on the assumptions of Section A and under two scenarios for the treatment effect: Trial 1 with an assumed hazard ratio of 1 and Trial 2 with an assumed hazard ratio of 0.75. We repeatedly simulated new trials until we had a trial that was stopped for fultility at interim analysis 1 (Trial 1) and one which was stopped for efficacy at interim analysis 2 (Trial 2). These simulated trials are used as the basis for suggested reporting in Sections B.1 and B.2.
When reporting results of a group-sequential design, or more general adaptive design, we strongly recommend to also follow the CONSORT extension for adaptive designs[7], as applicable.
B.1 Reporting of a trial that was stopped for futility
@@
B.2 Reporting of a trial that was stopped for efficacy